On December 21, 2020, you purchased 100 shares of ABC company at $11 per share. You plan to sell your shares on December 21, 2021 and are concerned about  downside risk. A put option on ABC stock with an exercise price (K) of $40 is currently priced (P) at $2 per share. Also, two call options on ABC stock with exercise prices (K) of $40 and $65 are priced (C) at $2.5 and $1.50 per share, respectively. All options expire on December 21, 2021.  What will be net profit/loss per share on a long bull call spread if the stock price is $70 per share?

A.

$14

B.

$24

C.

$34

D.

$44
Answer #1

Call spread: buy $40 strike call option and sell $65 strike call option.

Both options are exercised.

Profit on $40 call bought = 70 – 40 = 30

Loss on $65 call sold = 65 – 70 = -5

Net profit = 30 – 5 – 2.5 + 1.5 = $24

Answer is $24

Please rate.